speed of water in certain river is 6kmph and it takes a boat four times as long to row up as to row down the river. what is the speed of boat in still water?

Answer by subudear(62)

Let speed of water in river be u kmph
Let speed of boat in still water (water not moving) be v kmph
then time taken by boat to row DOWN the river will be
t*(v + u) = d ...(1)
t = time taken to row down, speed of boat will be (u+v) as flowing water will help boat go fast, d=distance traveled down stream
Similarly time time taken by boat to row UP the river will be
4t*(v - u) = d .....(2)
t = time taken to row UP which is 4 times the time taken to row DOWN, speed of boat will be (v - u) as flowing water will slowdown the boat, d=distance traveled down stream
As distance is same so equating both equations
t*(v + u) = 4t*(v - u)
This gives 5*u = 3*v
Now it is given that water speed is 6 kmph which means u = 6 kmph
substituting the value we get
v = 10 kmph
Speed of boat in still water is 10 kmph